OpenStudy (anonymous):

find the center of mass of a thin plate of a constant density d, covering the region bounded by the parabola y=5x^2 and the line y=20

OpenStudy (anonymous):

centre of mass will certainly lie on the y axis by symmetry we can say that, now to find its y coordinate, Ycm=integral(y.dm )/integral dm calculating the numerator, consider thin strip of width dy at a distance y on the parabola dm= density * area =D*x*dy now x=(y/5)^1/2 integral (y.dm) = integ{y*(y/5)^1/2*D*dy } D(1/5)^1/2integ{y^3/2.dy } D(1/5)^1/2*{ y^5/2}*2/5 now calculating integ dm ={D*x*dy} =D(1/5)^1/2integ{y^1/2.dy} =D(1/5)^1/2* y^3/2 *2/3 now dividing integral (y.dm) and integ dm we get 3/5y (good question)

OpenStudy (anonymous):

did u understand the solution?

OpenStudy (anonymous):

still digesting the info...i'm working on my calc 2 final and for whatever reason totally blanked out and am suffering from a brain fart so i'm cramming to get it done in an hour :(

OpenStudy (anonymous):

LOL, hmm thats pretty brave, best of luck

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

that previous question about finding volume that is some question, very good question, its ppl like you who make OS worth visiting

OpenStudy (anonymous):

SERIOUSLY